OpenStudy (anonymous):

Two vessels A and B of equal capacities contain mixtures of milk and water in the ratios 4: 1 and 3:1, respectively. 25% of the mixture from A is taken out and added to B. After mixing it thoroughly, an equal amount is taken out from B and added back to A. The ratio of milk to water in vessel A after the second operation is:

OpenStudy (anonymous):

5x-1/4-->1

OpenStudy (anonymous):

20x-1/4

OpenStudy (anonymous):

4x+(20x-1)/4-->2

OpenStudy (anonymous):

36x-1/4

OpenStudy (opcode):

Are there any answer choices?

OpenStudy (anonymous):

yes

OpenStudy (opcode):

Can you share them?

OpenStudy (anonymous):

79 : 21 83 : 1 77 : 23 81 : 19

OpenStudy (opcode):

Two vessels A and B of equal capacities contain mixtures of milk and water in the ratios 4: 1 and 3:1. Let's assume that each vessel is 100 litres: So for the 4:1 ratio it would be: Milk = 80 and water = 20 For the ration 3:1 it would be: Milk = 75 , water = 25 25% of the mixture from A is taken out and added to B. After mixing it thoroughly, an equal amount is taken out from B and added back to A. The ratio of milk to water in vessel A after the second operation is: 25 litres from A Would leave us remaining with 60 milk and water = 15 Vessel B would now have milk = 95 and water = 30 Ratio in B = 19:6 25 litres from B is transferred from B to A. So milk in A = 60 + 19 = 79 and water = 15 + 6 = 21 Ratio = 79:21